Systems Staff Meeting Notes 6/8/98

Systems Staff Meeting Notes
June 8, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Sharon Henry, Suzy Shaw, Leota Whitfield

Review Minutes of 6/1/98
  1. Re: state ordering of personal PC's - doesn't seem to be going through.
  2. May not save any money at all.
  3. There is a service provision after the purchase.
Announcements
  1. Re: Dale's secondary office on the third floor of East - we were asked to remove hanging wires, implications were that the room above John Ingram's needed to be cleared out and readied to be used as office space for Dale. Rob Roberts was under the impression that it was just to be a storage space for her boxes. Status of the space is still unclear.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. New shipment of memory received. Bill needs to go over the list with Lee before distribution. Can give out up to 30 sets - need to keep the rest.
  2. Have ordered 40 Cows - 12 with built-in zip drives. 6-8 of them will be for special purposes.
  3. Laptops & a server have been ordered from Gateway.
  4. After the public service meetings, we've determined that we won't have to buy too much for the restricted public work stations, although we may have to do more wiring, depending on where they're set up.
  5. 10 Front Pages have been ordered from the Hub. Problems may arise if people start using the extensions. NT stand-alones will need to be used as the web server, not Unix. Could cause security risks.
  6. Bill has received a fairly large number of CD ROM updates. They may not be mounted for 4-8 weeks. People will just have to be disappointed.
  7. 6 MP printer was delivered to Library Accounting. It is in place, not installed.
  8. Marcia needs to take delivery of the 5 printers & cables intended for Access Services when she returns.
  9. End of the Year Purchasing - there have been problems with people getting back with quotes. Standard end-of-the-year crash on expenditure, expect frantic calls.
  10. We've been getting conflicting reports on the status of printers - whether we will reuse them for staff or if they are needed for the public printing projects.
Liaisons
  1. Bill will post to the liaisons regarding the installation of time-sync. It will be optional for work stations.
Networking
  1. Hub was switched out at NERDC on Sunday evening. They've gone back to the original 1900 series hub.
  2. NERDC put on Novell filters - will immediately direct the routers to the Smathers server.
  3. Power hubs - problem left by UF Net. Even after the filters, there are some routers not responding correctly to the servers. There have been some reports about outside users temporarily losing CD access. NERDC is working to get rid of the power hubs. It is effecting IPX conductivity.
  4. SNA bump outs seem to have stopped. They are no longer occurring at the Circ Desk or with Cathy Vaughn.
  5. The autolog problem was an FCLA mistake, it is now fixed.
  6. Winston fixed the problem with the SQL server. Winston will be carrying his beeper with him while away at the seminar, in case such a problem arises again.
  7. Bill installed a service pack on NT 4.
  8. NERDC's 'herd of nerds' had several suggestions for us while they were over here last week. One was that the time-sync may be causing some of our problems .
  9. Facilities - John Humphrey had a printing problem, said Xerox had put some new software on Russ' machine. More than likely this was not Xerox, but the printing people. Xerox networking people won't be in until this Thursday morning.
  10. Switch in MSL is scheduled to take place on Thursday. The changeover shouldn't be a big deal, a walk through is in place for today.
  11. Will switch Music & AFA at the same time.
Software Projects Status
  1. ILL Lending has been migrated to the SQL server.
  2. Invoices ran fine.
  3. ILL Borrowing - Won't be migrated until July or August. Should be on the SQL server for this upcoming fiscal year.
  4. Winston is assisting Joan with end of the year work.
  5. OPS Tracking - networking.
  6. Sharon hopes to get something to Erika by next week, should be a major portion of her side.
  7. Brian - Paula will be down temporarily after the meeting for an upgrade.
  8. Brian has Special Collections about 90% finished, there's just 1-2 bugs to fix.
  9. Timer software for public work stations - need to get specs from Ashcraft.
  10. Sharon will be running NOTIS backups today.
  11. ATC standalone catalog - need to start thinking about getting back on that project - there will be interest in it statewide.
  12. Office'97 switch should be complete by now.
DP Coordinator
  1. Debra received a group of 12 test tapes from Compustat. They have changed the format of their yearly tapes & sent samples for testing purposes. Debra mapped the tapes & made them available for testing at NERDC. Peter McKay contacted finance professors in business, but they decline the opportunity to test the new format. When we receive the genuine updates in August, they will modify their programs to access the data.
  2. Debra needs to calculate the student budget for the fiscal year end.
Calendar
  1. Bobbie out June 8 - 19.
  2. Debra out half day June 16, all day June 17.
  3. Winston out June 15-24.
  4. ALA in Washington, D.C. is June 24 - 30.
  5. Rich will be gone to ALA/vacation June 24-July 15.
  6. Winston out July 2.
  7. Independence Day observed July 3.
  8. Brian and Lee will be out July 6.
  9. Debra out July 7 - 10.
  10. Bobbie out July 13 - 17
  11. Brian out July 16 & 17.
  12. Lee out Sept. 17-30.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu